# A Survey of Generative AI for Intelligent Transportation Systems

### v1

@@ 7.1. Open Challenges
Multi-modal traffic data. Multi-modal traffic data integrates various data sources and types, including roadside camera videos, traffic signal sensor data, weather conditions, traffic incident reports, and vehicle GPS trajectories. These data provides a comprehensive view of dynamic traffic conditions in transportation systems. Leveraging generative AI technology to address traffic-related problems with multi-modal data present several challenges. Firstly, the correlations of multi-modal data are more complex than those of single-modal data. For example, the characteristics of images and texts, as well as the differences in data distributions and representation methods, are distinct. Secondly, accurate alignment between different modalities is required, e.g., matching traffic camera images with traffic sensor data or weather information at the same time. Additionally, training models on multi-modal data can be more unstable, especially when maximizing mutual information across modalities. Therefore, how to effectively use generative AI technology to tackle transportation challenges using multi-modal data remains challenging in the field of transportation.
@@
Complex spatio-temporal correlations. Traffic involves complex relationships in time, space, and spatiotemporal interactions. In the temporal dimension, traffic data can be influenced by seasonal variations such as weekdays, weekends, and holidays. Simultaneously, the traffic flow at a specific moment can be impacted by the traffic patterns at that location over the preceding few hours. In the spatial dimension, traffic data involves the traffic conditions at different locations, and it is necessary to consider the traffic flow and congestion between different locations, and analyze their correlation. Moreover, there is a complex interaction between time and space in traffic conditions, and the traffic flow at a location may be affected by the traffic flow at adjacent locations in the past few hours. While prior research has explored the use of RNNs, CNNs, and GNNs in generative AI-related solutions to model spatiotemporal relationships, yielding some promising results, most of these approaches focus on local or micro-scale spatiotemporal modeling and fail to capture global or macro-scale spatiotemporal patterns. Furthermore, spatiotemporal modeling remains somewhat disconnected from the core principles of generative AI. Thus, the challenge in leveraging generative AI technology to address traffic-related issues lies in effectively modeling the intricate spatiotemporal correlations within traffic data.
@@
Sparse or missing traffic data. The sparsity or missing of traffic data can be attributed to various factors, including equipment failures or routine maintenance, network communication issues, and concerns related to the costs and privacy of data collection. The sparse or missing nature of traffic data can have significant implications for traffic management and planning, including inaccurate traffic predictions and an incomplete understanding of traffic conditions. Generative AI has powerful data generation capabilities and can generate relatively realistic data by leveraging the characteristics of existing data. To address the problem of data sparsity or absence, many researchers use generative AI technology to perform tasks such as data imputation, traffic data estimation, and traffic scene generation, but there are still shortcomings. First, due to the spatial-temporal complexity of traffic data, generative AI models may generate some seemingly realistic data that is inconsistent in space and time. Second, when generating data, the generative model may tend to generate common patterns in training data while neglecting less common cases. This may lead to a lack of diversity in the generated data and inability to cover all the changes in traffic data. Third, different degrees of sparsity also have a significant impact on the generation quality, and how to flexibly cope with this is an important issue. Therefore, overcoming the challenges of data sparsity and absence in addressing traffic issues with generative AI technology remains an important problem that needs further research.
@@
Adversarial attacks. In the realm of intelligent transportation, generative AI technology offers cutting-edge solutions that drive continuous advancements in traffic management and optimization. However, the widespread adoption of this technology also introduces new security risks, particularly those associated with adversarial attacks (Foo et al., 2023). One of the most notable concerns is the backdoor attack, where malicious actors may inject harmful data during the model training phase, causing the model to generate inappropriate outputs under specific conditions. In complex intelligent transportation networks, a single misjudgment or erroneous decision can trigger a chain reaction of traffic accidents or congestion, resulting in substantial economic and societal losses. For instance, an intelligent transportation monitoring system compromised by a backdoor attack may erroneously classify a safe vehicle as a potential threat, or an autonomous vehicle may make incorrect driving decisions as a result. Therefore, the critical challenge lies in accurately identifying potential risk points and devising effective methods to ensure the efficient and secure operation of transportation systems.
@@
Model interpretability. Model interpretability is crucial in intelligent transportation systems. For example, in autonomous driving scenarios, it is necessary to predict the behavior of other road users (pedestrians, cyclists, other vehicles) to ensure safe driving. This requires the model to explain why a particular behavior is considered the most likely, such as why the model believes a pedestrian in front may cross the road. Although generative AI models perform well in processing traffic data and optimizing transportation systems, these models are often considered black boxes, making it difficult to understand their decision-making processes and reasoning foundations. However, designing interpretable generative AI models is challenging. This is because traffic systems are very complex, involving many variables and uncertain factors such as road conditions, traffic flow, pedestrian behavior, and other vehicle behavior. These complexities and uncertainties make it difficult to design a simple and accurate model to explain traffic data and autonomous driving decisions. In autonomous driving, autonomous driving decisions involve multiple levels, including perception, planning, and control, and designing an interpretable model requires being able to explain the decision-making processes at these different levels, which are often very complex.
@@
Real-time requirements. In intelligent transportation systems, taking timely and effective response measures is of paramount importance. This becomes particularly critical when dealing with rapidly changing situations and unforeseen events. In such scenarios, traffic-related models must possess real-time perception capabilities, including the detection of other vehicles, pedestrians, and obstacles, while simultaneously making split-second decisions to ensure safe driving. In addition, traffic congestion can lead to rapid changes in traffic flow, so these models need to have the ability to predict real-time changes in traffic flow to take measures in advance and avoid traffic accidents and congestion. However, achieving this high level of responsiveness poses some challenges in generative AI-based solutions. First, real-time data may be affected by noise and incompleteness, so models must have the ability to handle imperfect data to accurately perceive the traffic environment. Second, certain generative AI methods may be very complex and require more time for reasoning. In real-time applications, it is necessary to balance the complexity and real-time performance of the model. Third, real-time applications should be equipped to handle exceptional circumstances like sensor failures or sudden events. Generative AI methods need to be capable of identifying such anomalies and executing appropriate actions to ensure both safety and reliability.
@@ 7.2. Potential Research Directions
Designing and training well-performed large-scale traffic models. Traditional transportation models and algorithms are typically developed separately for specific tasks, necessitating significant time and resource investments for model adjustments when encountering new tasks or changing problem requirements. Furthermore, these conventional models often exhibit excessive specialization, rendering cross-task transfer learning challenging and thereby limiting their overall generalizability and scalability. In contrast, large-scale language models based on generative AI technology have higher generalizability and scalability, and possess the potential for transfer learning in a variety of traffic tasks. For instance, in autonomous driving, which encompasses continuous tasks such as perception, prediction, and decision-making, we can design a large-scale transportation model to handle all three tasks simultaneously. In model design, special attention needs to be paid to shared features, such as spatiotemporal characteristics, and these features should be modeled effectively while combining the uniqueness of each task to achieve consistent excellent performance across all tasks. Additionally, another potential research direction is to develop large-scale traffic models driven by both data and knowledge, leveraging the diversity of traffic data and the reliability of traffic physical knowledge to effectively model traffic patterns. Finally, it is crucial to consider how to ensure that the designed large-scale model can achieve the desired performance and effectiveness. This requires the development of reasonable dataset annotation strategies and training strategies.
Traffic planning and decision-making based on large-scale language models. Traffic planning and decision-making are important research problems in the transportation domain. In real-world scenarios, traffic conditions are constantly changing, with the possibility of incidents such as traffic congestion, accidents, and adverse weather conditions, all of which may impact the effectiveness of traffic planning and decision-making. Therefore, it is essential to develop real-time traffic planning and decision-making algorithms based on large-scale language models. This entails exploring how to combine the real-time traffic perception capabilities of generative AI technology with the decision-making capabilities of RL. Such an approach holds the potential to yield more efficient and adaptable solutions to tackle the intricacies and uncertainties inherent in traffic management. As an illustrative example, consider driving decision-making in autonomous vehicles. Vehicles need to make decisions such as overtaking, decelerating, and stopping based on perceived data and environmental conditions. Large-scale models, equipped with powerful knowledge reasoning and generation capabilities, are expected to achieve precise traffic decisions in complex traffic environments. Therefore, future research can focus on how to design algorithms based on large-scale language models to make accurate decisions while ensuring a balance in real-time performance and safety.
@@ 8. Conclusions
Generative AI technology is playing an increasingly important role in perception, prediction, simulation, and control of transportation systems, providing powerful tools and methods to enhance traffic efficiency, safety, and sustainability. In this survey, we systematically investigate the critical solutions of generative AI techniques within intelligent transportation systems, and explore the challenging issues that generative AI has not yet fully addressed in this field. We hope that this survey will promote further innovation and development of generative AI technology in the field of intelligent transportation systems, assist researchers in effectively utilizing generative AI technology to tackle more practical traffic problems, and encourage the development of more advanced generative AI methods targeting the transportation domain.

### v2

@@ 7.1. Open Challenges
Multi-modal traffic data. Multi-modal traffic data combines various sources, such as roadside camera videos, traffic signals, weather conditions, incident reports, and vehicle GPS trajectories, providing a comprehensive view of traffic conditions. However, using generative AI to solve traffic-related problems with multi-modal data presents challenges. Firstly, the correlations between multi-modal data are more complex compared to single-modal data, as the characteristics and distributions of images, text, and other formats differ significantly. Secondly, accurate alignment between modalities, such as matching traffic camera images with sensor or weather data, is critical. Additionally, training models with multi-modal data can be unstable, especially when maximizing cross-modal information. Therefore, how to effectively use generative AI technology to tackle transportation challenges using multi-modal data remains challenging in the field of transportation.
@@
Complex spatio-temporal correlations. Traffic involves complex relationships across time, space, and spatiotemporal interactions. In the temporal dimension, traffic data is influenced by factors like weekdays, weekends, and holidays, and current traffic flow can be impacted by patterns from the preceding hours. In the spatial dimension, traffic conditions at different locations are interrelated, requiring analysis of correlations between traffic flow and congestion across locations. Additionally, traffic flow at one location may be affected by nearby traffic in the recent past. Previous research has employed RNNs, CNNs, and GNNs to model spatiotemporal relationships in traffic, showing promising results. However, most approaches focus on local or micro-scale spatiotemporal modeling and fail to capture global or macro-scale spatiotemporal patterns. Furthermore, spatiotemporal modeling has yet to fully utilize the powerful inference capabilities of generative AI. The main challenge lies in effectively modeling the intricate spatiotemporal correlations within traffic data to address traffic-related issues.
@@
Sparse or missing traffic data. Traffic data sparsity or gaps can arise due to equipment failures, maintenance, network issues, or the costs and privacy concerns of data collection. Such data gaps can lead to inaccurate traffic predictions and incomplete insights for traffic management. Generative AI, with its powerful data generation capabilities, has been used to address this issue through data imputation, traffic estimation, and scene generation. However, there are still shortcomings. First, due to the spatiotemporal complexity of traffic data, generative AI models may produce data that appears realistic but lacks spatial or temporal consistency. Second, when generating data, the generative model may tend to generate common patterns in training data while neglecting less common cases. This may lead to a lack of diversity in the generated data and inability to cover all the changes in traffic data. Third, different levels of data sparsity can significantly affect generation quality, requiring models to adapt flexibly. Therefore, overcoming the challenges of data sparsity and absence in addressing traffic issues with generative AI technology remains an important problem that needs further research.
@@
Adversarial attacks. In intelligent transportation, generative AI technology offers cutting-edge solutions for traffic management and optimization. However, its widespread adoption also introduces new security risks, particularly from adversarial attacks (Foo et al., 2023). One major concern is the backdoor attack, where malicious data injected during training can cause models to produce harmful outputs under specific conditions. In complex intelligent transportation networks, a single misjudgment malicious data injected during training can cause models to produce harmful outputs under specific conditions. For instance, an intelligent transportation monitoring system compromised by a backdoor attack may erroneously classify a safe vehicle as a potential threat, or an autonomous vehicle may make incorrect driving decisions as a result. Therefore, the critical challenge lies in accurately identifying potential risk points and devising effective methods to ensure the efficient and secure operation of transportation systems.
@@
Model interpretability. Model interpretability is crucial in ITS. For example, in autonomous driving scenarios, predicting the behavior of other road users, such as pedestrians, cyclists, and vehicles, is crucial for safe driving. This requires the model to explain why a particular behavior is considered the most likely, such as why a pedestrian may cross the road. While generative AI models excel in processing traffic data and optimizing transportation systems, they are often seen as black boxes, which makes it difficult to understand their decision-making. Designing interpretable generative AI models is challenging due to the complexity of traffic systems, involving numerous variables and uncertainties like road conditions, traffic flow, and pedestrian behavior. In autonomous driving, autonomous driving decisions involve multiple levels, including perception, planning, and control, and designing an interpretable model requires being able to explain the decision-making processes at these different levels, which are often very complex.
@@
Real-time requirements. In ITS, timely and effective response measures are crucial, especially in rapidly changing situations or unforeseen events. Traffic models must have real-time perception capabilities to detect vehicles, pedestrians, and obstacles while making immdiate decisions to ensure safety. Additionally, they must predict changes in traffic flow to take proactive measures against congestion and accidents. However, achieving this high level of responsiveness poses some challenges in generative AI-based solutions. First, real-time data may be affected by noise and incompleteness, so models must have the ability to handle imperfect data to accurately perceive the traffic environment. Second, certain generative AI methods may be very complex and require more time for reasoning. In real-time applications, it is necessary to balance the complexity and real-time performance of the model. Third, real-time applications should be equipped to handle exceptional circumstances like sensor failures or sudden events. Generative AI methods need to be capable of identifying such anomalies and executing appropriate actions to ensure both safety and reliability.
High-perfomance computing. In large metropolis, high population density results in diverse activities, complex travel behaviors, and numerous transportation routes, all of which present significant challenges for evaluating and optimizing traffic management in ITS (Aboudina et al., 2018). Further, leveraging generative AI for large-scale traffic simulations and optimizations involves running multiple simulations simultaneously, performing intensive analyses to find optimal solutions, and processing large volumes of data. High-Performance Computing (HPC) offers opportunities to provide the required computational power to meet these intensive demands. However, applying HPC to support generative AI in ITS presents several challenges. As the massive amounts of data generated from numerous sources in ITS applications, highly efficient data handling and parallel processing are crucial for enabling timely decision-making. Another key challenge is ensuring scalability and efficiency in real-time applications. ITS environments are highly dynamic, with traffic conditions changing rapidly due to factors such as accidents, weather changes, or special events. The HPC framework must be able to scale computational resources to handle sudden spikes in data volume or computational demand, ensuring the system remains responsive under varying conditions.
@@ 7.2. Potential Research Directions
Designing and training well-performed large-scale traffic models. Traditional transportation models are typically developed separately for specific tasks, requiring significant time and resources to adapt to new requirements. These models also tend to be overly specialized, making cross-task transfer learning difficult and limiting scalability. In contrast, large-scale language models based on generative AI technology have higher generalizability and scalability, with the potential for transfer learning across various traffic tasks. For instance, in autonomous driving, which encompasses continuous tasks such as perception, prediction, and decision-making, a large-scale model could handle all tasks simultaneously by modeling shared features like spatiotemporal characteristics while considering each task’s unique aspects. Recently, GPT-ST (Li et al., 2023d) gives a preliminary attempt in designing generative spatio-temporal pre-training framework. Another promising direction is developing large-scale traffic models driven by both data and knowledge, combining diverse traffic data with reliable physical knowledge to effectively model traffic patterns.s
Traffic planning and decision-making based on large-scale language models. Traffic planning and decision-making are important research problems in the transportation domain. In real-world scenarios, traffic conditions constantly change due to congestion, accidents, or adverse weather, impacting the effectiveness of planning and decision-making. Therefore, it is essential to develop real-time traffic planning and decision-making algorithms based on large-scale language models. This entails exploring how to combine the real-time traffic perception capabilities of generative AI technology with the decision-making capabilities of RL. Such an approach holds the potential to yield more efficient and adaptable solutions to tackle the intricacies and uncertainties inherent in traffic management. As an illustrative example, consider driving decision-making in autonomous vehicles. Vehicles need to make decisions such as overtaking, decelerating, and stopping based on perceived data and environmental conditions. Large-scale models, equipped with powerful knowledge reasoning and generation capabilities, are expected to achieve precise traffic decisions in complex traffic environments. Therefore, future research can focus on how to design algorithms based on large-scale language models to make accurate decisions while ensuring a balance in real-time performance and safety.
@@ 8. Conclusions
Generative AI technology is playing an increasingly influential role in ITS, particularly in perception, prediction, simulation, and decision-making. In this survey, we systematically investigate key generative AI solutions within ITS, and explore the challenges that remain unresolved in this field. Our analysis identifies several valuable guidelines. First, designing specialized HPC allows for large-scale traffic simulations and optimizations with massive multi-modal data. Second, ensuring data reliability and system security requires addressing data sparsity and safeguarding against adversarial attacks. Third, enhancing real-time performance while improving model interpretability is crucial to enable prompt responses to dynamic traffic conditions and foster trust through transparent decision-making. Based on these guidelines, we summarize several key trends in generative AI for ITS. One significant trend is the integration of multi-modal data to enable comprehensive and large-scale traffic modeling. Another is the focus on designing generalizable models capable of addressing multiple transportation tasks simultaneously, which enhances the versatility and robustness of AI solutions. Additionally, there is growing interest in developing real-time traffic planning and decision-making algorithms, particularly by leveraging the capabilities of large-scale language models. These trends illustrate the transformative potential of generative AI in reshaping ITS, paving the way for safer, more efficient, and resilient transportation systems.